Alignment of the H1 OMC cavity mirrors
- The cavity mirrors as well as the first steering mirror were aligned on the cavity side template.
- The locking of the cavity was not so stable as before. Some high freq (several hundreds Hz) disturbance makes the cavity
deviate from the linear range. This can be mitigated by turning off the HEPA units but this is not an ideal condition.
- FSR and TMS were measured.
These suggest the cavity length L and f_TMS/f_FSR (say gamma, = gouy phase / (2 pi) ) as
L=1.1343 m (1.132m nominal)
gamma_V = 0.219659 (0.21879 nominal)
gamma_H = 0.220484 (0.21939 nominal)
- the 9th modes of the carrier is away from the resonance 6-9 times of the line width (LW)
- the 13th modes of the lower f2 sideband are 11-15 LW away
- the 19th modes of the upper f2 sideband are 0.6-7 LW away
We still need precise adjustment of the gouy phase / cavity length, this was enough for the gluing of the flat mirrors